OpenStudy (anonymous):

The ratio of the number of apples to the number of oranges is 5:7. The ratio of the number of oranges to bananas is 8:9. There is a total of 3180 fruits. How many bananas are there? Please help!

OpenStudy (anonymous):

mutliply 8*9

OpenStudy (retireed):

Let A = apples, OR = oranges and B = Banana Write the ratio in proportional form (see diagram) Cross multiple, solve in terms of A = ??? OR and B = ??? OR Then substitute A and B into the sum of fruits equation A + OR + B =3180 and solve for OR (the math can get a bit tricky) Once you know OR, you can solve for B with equation B = ??? OR 

OpenStudy (retireed):

Eq 1 A = (5/7) OR Eq 2 B = (9/8) OR Eq 3 A + OR + B = 3180 Eq 4 (5/7) OR + OR + (9/8) OR = 3180 OR = ??? so now B = (9/8) times ???

OpenStudy (anonymous):

Thanks for your help...........that was easy to get with your help!